OpenStudy (anonymous):

Give exact and approximate solutions to three decimal places? (x+9)^2=4

OpenStudy (anonymous):

\[(x+9)^2=4\] \[x+9=2\] or \[x+9=-2\] \[x=-7\] or \[x=-11\]